OpenStudy (anonymous):

Palmer's average running speed is 3km/h faster than his walking speed. If Palmer can run around a 30-kilometer course in 2h, how many hours would it take for Palmer to walk the same course?

OpenStudy (anonymous):

30km/2h would give you his running speed. So the answer to that minus 3 would give you his walking speed. Divide 30km by that answer to get the number of hours.

OpenStudy (anonymous):

30/2=15-3=12 12/30= the answer?

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

noooo 30/12.

OpenStudy (anonymous):

" running speed is 3km/h faster than his walking speed" Let r = running speed, w is walking speed -->r = w + 3 ( km/h ) Agree?

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

" run around a 30-kilometer course in 2h" -> r = 30 km/ 2 hr = ...?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

15+3?

OpenStudy (anonymous):

W=15?

OpenStudy (anonymous):

w = r - 3

OpenStudy (anonymous):

r = 15 => w = ..?

OpenStudy (anonymous):

time = distance / walking speed = 20 / 12 = ... hr

OpenStudy (anonymous):

Where does 20 comes from?

OpenStudy (anonymous):

Oops 30/12 =

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

= 2 hr 30 min

OpenStudy (anonymous):

Such a saviour ;-)

OpenStudy (anonymous):

I hope you can learn from it ;)

OpenStudy (anonymous):

I will gonna study and understand this for my exam Thanks a lot;-)